# H2R: A Human-to-Robot Data Augmentation for Robot Pre-training from Videos

### v3

@@ Abstract
Large-scale pre-training using videos has proven effective for robot learning. However, the models pre-trained on such data can be suboptimal for robot learning due to the significant visual gap between human hands and those of different robots. To remedy this, we propose H2R, a simple data augmentation technique that detects human hand keypoints, synthesizes robot motions in simulation, and composites rendered robots into egocentric videos. This process explicitly bridges the visual gap between human and robot embodiments during pre-training. We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2, replacing human hands with simulated robotic arms to generate robot-centric training data. To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions. We validate H2R across four simulation benchmarks—Robomimic, RLBench, PushT and CortexBench, and real-world manipulation tasks with UR5 and Dual-Arm Franka/UR5. H2R consistently improves downstream success rates, yielding gains of 1.3%–10.2% in simulation and 3.3%–23.3% in real-world tasks across various visual encoders and policy learning methods. We further demonstrate the potential of H2R in cross-embodiment generalization and its compatibility with VLA models. These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains. The project page is available in https://sites.google.com/view/h2r-robotics.
1 Introduction
Pre-training of generalizable robotic features for object manipulation and motion navigation constitutes a crucial objective within the realm of robotics. Inspired by the remarkable accomplishments of large scale pre-training in computer vision [he2022mae, nair2022r3m, zeng2024learningmanipulationpredictinginteraction, karamcheti2023voltron, datacomp] and natural language processing [devlin2018bert, yu2022pointbert, openai2024gpt4technicalreport, anil2023palm2technicalreport, li2024datacomplm], many efforts have been devoted to harnessing large-scale data to construct generalizable representations in the robotics field [embodimentcollaboration2024openxembodimentroboticlearning, brohan2023rt1roboticstransformerrealworld, brohan2023rt2visionlanguageactionmodelstransfer]. Nevertheless, when it comes to robot manipulation, the process of collecting demonstrations is labor-intensive and expensive [duan2023ar2, jang2022bczzeroshottaskgeneralization, brohan2023rt1roboticstransformerrealworld, khazatsky2024droidlargescaleinthewildrobot, fang2023rh20tcomprehensiveroboticdataset, gao2024efficientdatacollectionrobotic, lee2021pickandplacetacklingroboticstacking, herzog2023deeprlscalesorting, kalashnikov2021mtoptcontinuousmultitaskrobotic, duan2024manipulate]; meanwhile, there exist many large-scale egocentric video datasets showing how humans perform manipulation and navigation, which can potentially serve as a cheap alternative of demonstrations for the pre-training of generalizable visual features for robotics.
Recent works [nair2022r3m, xiao2022mvp, zeng2024learning] analyze such egocentric human video datasets such as Ego4D [grauman2022ego4d], SSv2 [goyal2017something], and Epic Kitchens [damen2018scaling] with the aim of gleaning prior knowledge about object manipulation and enabling the acquisition of general and robust feature representations. However, during the representation learning, the gap in visual representations between the human arm and the robotic arm remains largely unaddressed and can hinder the transferability of models trained on egocentric datasets to robotic systems. Specifically, when utilizing the robot expert data to fine-tune the pre-trained robotic representations for downstream robotic tasks, the model has to learn to bridge the visual gap between the first-person human hand and the robots in addition to acquiring nuanced task-specific skills demonstrated in the robot expert data. This would result in increased complexity during the fine-tuning process and suboptimal performance.
@@ 1 Introduction
To mitigate this issue, we propose H2R (as shown in Figure LABEL:fig:pipeline), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation. H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video. Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR [pavlakos2023reconstructinghands3dtransformers] to accurately detect the position and posture of the human hand in egocentric videos. Then, we simulate the same robot state in simulators to obtain the mask of robot actions. In the second stage, we use the Segment Anything Model [kirillov2023segment] to automatically separate human hand from background, and use the inpainting model LaMa [suvorov2021resolutionrobustlargemaskinpainting] to fill the removed hand mask. After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video. Finally, we overlay the robotic arm images captured by the simulator’s camera onto the areas where the human hands are removed. Through such a process, H2R explicitly reduces the gap between human and robot hands by creating realistic robotic arm movements that visually mimic human hand actions. It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
@@
To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics. This provides a lightweight and scalable proxy to assess the alignment quality between input human videos and robot-augmented outputs.
To further verify the utility of the augmented datasets, we conduct downstream experiments comparing models pre-trained on original egocentric datasets with those trained on our released H2R-enhanced datasets. These comparisons are performed on both simulation and real-world robotic manipulation tasks to assess how H2R impacts policy learning performance in practice. For the pre-training stage, we apply MAE [he2022mae] and R3M [nair2022r3m] frameworks to train visual encoders using 62,500 videos from the SSv2 dataset, where 16 keyframes are sampled per video. The pretrained visual representations are used in downstream imitation learning pipelines. We conduct comprehensive evaluations across both simulation and real-world manipulation tasks. In simulation, we evaluate pre-trained visual encoders on four benchmark suites: Robomimic, RLBench, PushT and CortexBench. H2R improves the average success rates by +10.2% for MAE and +6.3% for R3M on Robomimic tasks, +10.0% for MAE and +5.0% for R3M on RLBench tasks, +5.3% for MAE and +7.0% for R3M on the PushT task, and +5.3% for MAE and +1.3% for R3M on the CortexBench task. In real-world experiments, we deploy H2R-enhanced encoders on different robot and end-effector settings, evaluated over nine manipulation tasks under two policy frameworks: Diffusion Policy (DP) [chi2023diffusionpolicy] and ACT [zhao2023act]. H2R brings consistent improvements, with average success rate gains of +10.0% (MAE) and +14.1% (R3M) under DP, and +12.8% (MAE) and +13.3% (R3M) under ACT. H2R also improves MAE/R3M performance when pre-trained on Ego4D, with up to +10% and +15% gains in simulated and real-world tasks, respectively. Beyond these evaluations, we also test H2R’s ability to support cross-embodiment transfer by pre-training encoders on Franka data and deploying them on UR robots, indicating that H2R improves representation transfer across different morphologies. In addition, fine-tuning VLA backbones on H2R-generated datasets yields markedly higher success rates on dual-arm UR tasks than fine-tuning on human-hand videos or using no fine-tuning. These results demonstrate that H2R effectively bridges the human-to-robot visual domain gap and significantly enhances downstream performance across both simulation and real-world manipulation settings.
2 Related Work
@@
Robot Imitation Learning. Data-driven policy learning [sergey2018learning, luo2024hilserl, xiao2022mvp, yang2024equibotsim3equivariantdiffusionpolicy, chi2023diffusionpolicy, embodimentcollaboration2024openxembodimentroboticlearning] has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation. Imitation learning [chi2023diffusionpolicy, ze2024dp3, zhao2023act, mandlekar2021matters] trains policies from successful demonstrations, often supervised by behavior cloning [torabi2018behavioral, florence2021implicit] objectives. ACT [zhao2023act] addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models [chi2023diffusionpolicy, ze2024dp3] are introduced to handle the inherent multimodality of robot motions. In addition, CordViP [fu2025cordvip] leverages 3D object-robot correspondences to enhance dexterous manipulation.
@@
Visual Encoder Pretraining for Robotics. Visual pretraining improves generalization of robotic policies across diverse tasks. Researchers have explored architectural designs [He_2016_CVPR, dosovitskiy12021an], training objectives [He_2022_CVPR, chen2020simple, chen2020big], and dataset compositions [deng2009imagenet, lin2014microsoft, shao2019object365, Schuhmann2021LAION400MOD]. PVR-Control [parisi22a] shows that pretrained visual representations can outperform direct state-based policies. RPT [radosavovic2023robot] tokenizes observations to enable masked prediction pretraining. Methods like MVP [radosavovic2022realworld] and R3M [nair2022r3m] utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning. Voltron [karamcheti2023voltron] demonstrates the use of MAE and contrastive learning for hierarchical robot control.
@@
Cross-Domain Visual Alignment. Bridging the domain gap between human and robot visual inputs remains a major challenge. WHIRL [bahl2022humantorobotimitationwild] matches task structure from third-person views, while RoVi-Aug [chen2024roviaugrobotviewpointaugmentation] and Mirage [chen2024mirage] manipulate appearance via segmentation or image-space preprocessing. EgoMimic [kareer2024egomimicscalingimitationlearning] removes hands and normalizes views to align egocentric perspectives.
3 H2R: Human-to-Robot Data Augmentation
In this section, we describe H2R, a data augmentation pipeline for robot learning from egocentric human videos (Figure 1). It replaces human hands in every frame with robotic arms equipped with various end effectors, generating a new, visually different dataset. This approach aims to mitigate the visual gap between human hands and robots, thereby improving the generalizability of visual representations learned from egocentric data to robotic domains. We first examine the position and pose of the hand and arm in the image, then map this pose to the robotic arm and end-effector in the simulator. Next, we erase the original hand and arm from the image and overlay a rendered image of the simulated robot performing the same pose onto the original image in place of the removed hand and arm. Figure 3 illustrates examples of human hand videos that have been processed using H2R.
3.1 H2R Data Augmentation Pipeline
Figure 1: H2R Pipeline. H2R involves replacing human hands with robotic arms by first using the HaMeR model to detect hand poses and camera parameters. The human hand is then removed using the SAM, and the inpainting model LaMa fills in the gap. A robot hand is constructed based on the detected pose and keypoints, with the camera perspective adjusted to match the original image. Finally, the robot hand is overlaid onto the image, ensuring accurate alignment with the human hand.
@@ 3.1 H2R Data Augmentation Pipeline
3D Hand Pose Estimation. In order to overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information. We adopt HaMeR [pavlakos2023reconstructinghands3dtransformers], a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image, providing precise positional information for subsequent hand removal. Given an ego-centric RGB image, HaMeR estimates both the hand pose (including 3D keypoints) and the intrinsic and extrinsic parameters of the rendering camera.
@@
Robotic Arm and End Effector Construction. This step involves building the robotic arm and its end-effectors, which are typically either grippers or dexterous hands. For dexterous hands, we estimate joint angles from hand keypoints predicted by HaMeR. Each finger joint angle is calculated using three consecutive keypoints along that finger. For grippers, we determine how open or closed they are based on the Euclidean distance between the corresponding fingertips. Since hand keypoints don’t capture the full arm pose (especially for joints unrelated to the hand), we manually set reasonable values for the remaining arm joints to complete a plausible robot configuration.
@@
Simulator Camera Position Alignment. The visual bias introduced by the camera perspective is more significant than the action retargeting itself. To address this, we use the hand keypoints and camera parameters from HaMeR to adjust the camera pose in the simulator. Specifically, we define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robotic arm in the simulator. By mapping the position of the camera in $C_{H}$ to $C_{S}$, we can ensure that the camera in the simulator shares the same perspective as the one captured in the real-world egocentric human image. The original camera position $^{W}\mathbf{cam}_{real}$ in the world frame is transformed to the aligned simulator position $^{W}\mathbf{cam}_{sim}$ using transformations from human hand ($^{W}_{H}\mathbf{R}$) and robot simulator ($^{W}_{S}\mathbf{R}$) coordinate systems:
$^{W}\mathbf{cam}_{sim}=^{W}_{S}\mathbf{R}\times^{W}_{H}\mathbf{R}^{-1}\times^{W}\mathbf{cam}_{Real}$ (1)
A more detailed explanation of the coordinate transformation and camera alignment process is provided in Appendix.
Robot Hand Rendering and Copy-paste. After setting the camera, the segmentation mask of the robotic arm is obtained by rendering with the simulator camera. We directly obtain the pixel coordinates of the human hand keypoints from HaMeR, which are predicted from the input image. In parallel, the pixel coordinates of the robot end-effector links are computed in the simulator by projecting their 3D positions through the aligned camera using the known transformation matrices. By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.
@@
Human Arm and Hand Removal. After obtaining the hand pose information, we need to mask out the human hand and then inpaint the masked area with a proper background. We use Segment Anything Model (SAM) [kirillov2023segment] to automatically segment the human hand and arm regions using the hand pose information detected by HaMeR. Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa [suvorov2021resolutionrobustlargemaskinpainting], a powerful inpainting model, to fill in the removed hand-arm region. This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
3.2 Data Quality Evaluation
Figure 2: H2R samples. Visual comparison between original human data (top) and our augmented data (bottom).
To assess the visual plausibility and semantic consistency of the robot-augmented images produced by H2R, we employ a vision-language similarity evaluation based on CLIP [radford2021learningtransferablevisualmodels]. This method measures how well the rendered robot actions align with high-level semantic descriptions of manipulation tasks. For each augmented frame, we formulate a pair of textual prompts describing the same action: one from a human-centric perspective and one from a robot-centric perspective. The human-centric template is defined as: A human is [action]; the robot-centric template is defined as: A robotic arm is [action]. Here, [action] is a natural language phrase describing the high-level behavior, such as ”holding a bottle” or ”brushing paint on a wall”.
We use the CLIP ViT-B/32 model to compute the cosine similarity between the image embedding of the augmented frame and the corresponding textual prompt. This yields a scalar score reflecting how semantically consistent the image content is with the robot-centric description. Similarly, we compute the similarity between the original human frame and the human-centric prompt. This setup allows us to directly compare semantic alignment before and after augmentation under their respective modalities. For each frame, only the action-relevant region is retained, and background context is preserved via inpainting using LaMa. We average similarity scores over multiple samples per action class to mitigate variance and isolate the impact of robotic replacement. All textual prompts are manually curated to maintain consistency across frames and embodiments.
Figure 2 contains six pairs of subfigures, each showing a human action and its corresponding augmented robotic action. The six action pairs in Figure 2 (left to right) are described as follows: (1) Holding a yellow measuring tape at a construction site; (2) Holding a bottle; (3) Painting on the wall with a brush; (4) Rushing paint on a wall; (5) Turning on the kitchen faucet; (6) Preparing painting materials on a table.
The CLIP similarity scores are reported in Table 1. Across all samples, H2R-augmented images consistently achieve higher similarity with robot-centric action descriptions than original images do with human-centric ones. This confirms that H2R augmentation effectively improves visual-semantic alignment between human demonstrations and robotic execution.
Table 1: CLIP similarity scores. Higher values indicate better alignment between images and action descriptions.
Img-1 Img-2 Img-3 Img-4 Img-5 Img-6
Ori 30.6 23.7 31.2 31.7 27.5 28.7
Aug 32.6 28.9 32.3 32.0 28.7 31.0
This evaluation provides a lightweight, scalable, and interpretable metric for assessing how well visual hand-to-robot transformations preserve task semantics. To quantify data quality, we sample 1,000 image pairs (original and H2R-augmented). For each original image, we use Qwen-2.5-VL [bai2025qwen25vltechnicalreport] to generate a verb-noun action description (e.g., “washing bowl”). We then form two prompts: “A human is [action]” and “A robot arm is [action]”. Using CLIP ViT-B/32, we compute the cosine similarity between: (1) the human prompt and the original image, and (2) the robot prompt and the H2R-augmented image. We report the average similarity over all 1,000 samples. The raw images achieve an average score of 28.01, while H2R-augmented images reach 29.83. This demonstrates that H2R-augmented images exhibit stronger semantic alignment with robot-centric action descriptions, confirming improved visual-semantic coherence for robotic manipulation. The results validate our data curation pipeline as effective for generating high-quality, robot-relevant visual data at scale.
Figure 3: Examples of H2R Augmentation. Each column shows images before and after augmentation. The top row uses UR5 Leaphand, the middle row uses UR5 Gripper, and the bottom row uses Franka to replace human hands.
4 Simulation Experiment
Encoder Pre-training. We adopt the MAE [he2022mae, xiao2022mvp] and R3M [nair2022r3m] frameworks for pre-training, each employing a Vision Transformer (ViT) Base [dosovitskiy2021imageworth16x16words] model as the visual encoder. The training dataset used is a subset of Ego4D (117K clips) or SSv2 (1M images). Both MAE and R3M are pre-trained with $\timesA8$00 GPUs. MAE employs 800 epochs with 128 batch size and 4e-4 learning rate, while R3M uses 20K steps with 256 batch size and 1e-4 learning rate.
Policy Training. We train for 200 steps and report the mean success rate for Robomimic tasks. For tasks in RLBench [james2019rlbench], we train for 800 epochs for each policy model and test them 20 times in the RLBench environment with a random initialization. For the PushT task, we train the Diffusion Policy model for 200 epochs and report the success rate in the simulation environment. For VC-1 Cortex benchmark, we train policies using behavior cloning composed of a 3-layer MLP for 100 epochs. The training hyperparameters used in this work are identical to those described in the original paper.

### v4

@@ Abstract
Large-scale pre-training using egocentric human videos has proven effective for robot learning. However, the models pre-trained on such data can be suboptimal for robot learning due to the significant visual gap between human hands and those of different robots. To remedy this, we propose H2R, a human-to-robot data augmentation pipeline that converts egocentric human videos into robot-centric visual data. H2R estimates human hand pose from videos, retargets the motion to simulated robotic arms, removes human limbs via segmentation and inpainting, and composites rendered robot embodiments into the original frames with camera-aligned geometry. This process explicitly bridges the visual gap between human and robot embodiments during pre-training. We apply H2R to augment large-scale egocentric human video datasets such as Ego4D and SSv2. To verify the effectiveness of the augmentation pipeline, we introduce a CLIP-based image-text similarity metric that quantitatively evaluates the semantic fidelity of robot-rendered frames to the original human actions. We evaluate H2R through comprehensive experiments in both simulation and real-world settings. In simulation, H2R consistently improves downstream success rates across four benchmark suites—Robomimic, RLBench, PushT, and CortexBench—yielding gains of 1.3%–10.2% across different visual encoders and policy learning methods. In real-world experiments, H2R improves performance on UR5 and dual-arm Franka/UR5 manipulation platforms, achieving 3.3%–23.3% success rate gains across gripper-based, dexterous, and bimanual tasks. We further demonstrate the potential of H2R in cross-embodiment generalization and its compatibility with vision–language–action models. These results indicate that H2R improves the generalization ability of robotic policies by mitigating the visual discrepancies between human and robot domains.
I Introduction
Pre-training generalizable visual representations is a central challenge in robotic manipulation. Recent advances in large-scale pre-training in computer vision and language have significantly improved representation learning across domains [22, 37, 27, 17, 10, 1]. However, collecting large-scale robot demonstrations remains labor-intensive and costly [13, 29, 14], motivating the use of readily available egocentric human videos as an alternative source for robot pre-training.
Large-scale egocentric datasets such as Ego4D [20], Something-Something V2 [19], and EPIC-Kitchens [8] capture diverse human-object interactions and have been shown to support transferable visual representations for robotic manipulation. However, these datasets are inherently human-centric. The visual mismatch between human hands in egocentric videos and robotic embodiments at deployment time introduces a gap that is not explicitly addressed during pre-training, limiting representation transfer.
@@ I Introduction
To mitigate this issue, we propose H2R (as shown in Figure LABEL:fig:pipeline), a simple data augmentation method that converts videos of Human hand operations into that of Robotic arm manipulation. H2R consists of two major procedures: the first part is to generate the robotic movements to imitate the human hand movements in a video, followed by the second stage that overlays the robotic movements onto the human hand’s movements in the video. Specifically, in the first part, we employ state-of-the-art 3D hand reconstruction model HaMeR [40] to accurately detect the position and posture of the human hand in egocentric videos. Then, we simulate the same robot state in simulators to obtain the mask of robot actions. In the second stage, we use the Segment Anything Model [30] to automatically separate human hand from background, and use the inpainting model LaMa [51] to fill the removed hand mask. After that, we align the camera intrinsic parameters of the images detected in HaMeR with those in the simulator, and then achieve pixel-level matching between the robotic arm images in the simulators and the human hand images in the egocentric video. Finally, we overlay the robotic arm images captured by the simulator’s camera onto the areas where the human hands are removed. Through such a process, H2R explicitly reduces the gap between human and robot hands by creating realistic robotic arm movements that visually mimic human hand actions. It allows the model to learn the task-specific actions demonstrated by the human hand, but with robotic arm visual representations that are more suitable for robotic systems.
@@
To evaluate the effectiveness of the H2R augmentation process, we introduce a CLIP-based [41] semantic similarity metric that measures how well the rendered robot frames preserve the original action semantics. This provides a lightweight and scalable proxy to assess the alignment quality between input human videos and robot-augmented outputs.
We further evaluate the effectiveness of H2R through a comprehensive set of experiments in both simulation and real-world settings. In simulation, we conduct imitation learning experiments on four benchmark suites—Robomimic [36], RLBench [24], PushT [7], and CortexBench [35]— using visual encoders pre-trained with MAE and R3M, and downstream policies trained under standard behavior cloning and diffusion-based policy learning frameworks. Across these benchmarks, pre-training with H2R consistently improves downstream success rates, with average gains ranging from 1.3% to 10.2% depending on the encoder and task suite.
In real-world experiments, we deploy H2R-enhanced visual encoders on multiple robotic platforms, including a UR5 arm with parallel grippers, a UR5 arm with a dexterous Leaphand, and dual-arm systems based on Franka and UR5e. Policies are trained using both Diffusion Policy [7] and ACT [57]. Across gripper-based, dexterous, and bimanual manipulation tasks, H2R yields consistent performance improvements in real-world success rates, with gains ranging from 3.3% to 23.3% across different encoder–policy combinations.
Beyond these primary evaluations, we conduct additional studies to characterize the generalization properties of H2R. These include experiments on pre-training with different egocentric datasets, cross-embodiment transfer where the augmentation embodiment differs from the downstream robot, robustness analyses under varying demonstration densities and lighting conditions, and integration with vision–language–action models through targeted fine-tuning. Collectively, these results demonstrate that explicitly reducing the human-to-robot visual gap during pre-training leads to robust and transferable improvements in downstream robotic manipulation performance.
Figure 1: H2R Pipeline. The H2R pipeline includes three steps: (1) hand pose estimation, (2) retargeting to simulate robot arm movements, and (3) removal and inpainting of the human hand to create an augmented image with the robot arm, ensuring seamless integration.
II Related Work
@@
Robot Imitation Learning. Data-driven policy learning [31, 34, 53, 55, 7, 38] has enabled robots to autonomously perform tasks such as grasping, locomotion, and manipulation. Imitation learning [7, 56, 57, 36] trains policies from successful demonstrations, often supervised by behavior cloning [52, 15] objectives. ACT [57] addresses non-Markovian dynamics by fusing temporal sequences, while diffusion models [7, 56] are introduced to handle the inherent multimodality of robot motions.
@@
Visual Encoder Pretraining for Robotics. Visual pretraining improves generalization of robotic policies across diverse tasks. Researchers have explored architectural designs [23, 12], training objectives [21, 5, 6], and dataset compositions [9, 33, 48, 47]. PVR-Control [39] shows that pretrained visual representations can outperform direct state-based policies. RPT [44] tokenizes observations to enable masked prediction pretraining. Methods like MVP [45] and R3M [37] utilize self-supervised objectives on videos to learn representations transferable to reinforcement learning. Voltron [27] demonstrates the use of MAE and contrastive learning for hierarchical robot control.
@@
Cross-Domain Visual Alignment. Bridging the domain gap between human and robot visual inputs remains a major challenge. WHIRL [2] matches task structure from third-person views, while RoVi-Aug [4] and Mirage [43] manipulate appearance via segmentation or image-space preprocessing. EgoMimic [28] removes hands and normalizes views to align egocentric perspectives.
III H2R: Human-to-Robot Data Augmentation
In this section, we present H2R, a human-to-robot data augmentation framework that converts egocentric human-hand videos into robot-centric visual observations. We detail the full augmentation pipeline, including (1)hand pose estimation, (2)retargeting, and (3)removal and inpainting, and further validate the quality of the generated data using a CLIP-based semantic consistency evaluation. An overview of the pipeline is shown in Figure 1.
III-A H2R Data Augmentation Pipeline
@@
Step1: Hand Pose Estimation. To overlay the human hands in the egocentric image with different robots, we first need an efficient and accurate model to detect hand information. We adopt HaMeR [40], a state-of-the-art model for 3D hand detection and reconstruction, to accurately locate the hand in the image. This step outputs the position of the human hand, keypoints, and the intrinsic and extrinsic parameters of the camera. The hand position will be used in Step 3 (Removal and Inpainting) to segment the human hand, while the keypoints and camera parameters are used in Step 2 (Retargeting) to adjust the robot’s pose and camera position.
Step2: Retargeting. In this step, we simulate the robot arm to mirror the human hand’s movements detected in Step 1. Using the hand keypoints and camera parameters from Step 1, we calculate the joint angles of the robot arm.
@@
We first build the robotic arm and its end-effectors, which are typically either grippers or dexterous hands. For dexterous hands, we estimate joint angles from hand keypoints predicted by HaMeR. Each finger joint angle is calculated using three consecutive keypoints along that finger. For grippers, we determine how open or closed they are based on the Euclidean distance between the corresponding fingertips. Since hand keypoints don’t capture the full arm pose (especially for joints unrelated to the hand), we manually set reasonable values for the remaining arm joints to complete a plausible robot configuration.
@@
Next, we use the hand keypoints and camera parameters from HaMeR to adjust the camera pose in the simulator. Specifically, we define two coordinate systems: $C_{H}$, the coordinate system aligned with the human hand, and $C_{S}$, the coordinate system of the robotic arm in the simulator. By mapping the position of the camera in $C_{H}$ to $C_{S}$, we can ensure that the camera in the simulator shares the same perspective as the one captured in the real-world egocentric human image. The original camera position $^{W}\mathbf{cam}_{real}$ in the world frame is transformed to the aligned simulator position $^{W}\mathbf{cam}_{sim}$ using transformations from human hand ($^{W}_{H}\mathbf{R}$) and robot simulator ($^{W}_{S}\mathbf{R}$) coordinate systems:
$^{W}\mathbf{cam}_{sim}=^{W}_{S}\mathbf{R}\;^{W}_{H}\mathbf{R}^{-1}\;{}^{W}\mathbf{cam}_{real}$ (1)
Step3: Removal and Inpainting. After the robot pose is retargeted in Step 2, we proceed to remove the human hand from the image. We use the hand position and keypoints from Step 1 to segment out the human hand and arm regions using the Segment Anything Model (SAM) [30].
@@
Then, to obtain clean backgrounds for inserting robotic arms, we apply LaMa [51], a powerful inpainting model, to fill in the removed hand-arm region. This yields clean RGB images without human limbs, providing a seamless background for inserting robotic arms in the subsequent steps.
Once both the arm segmentation from Step 2 and the inpainted image from Step 3 are available, we overlay the robot arm onto the inpainted image. We directly obtain the pixel coordinates of the human hand keypoints from Step 1. In parallel, the pixel coordinates of the robot end-effector links are computed in the simulator by projecting their 3D positions through the aligned camera using the known transformation matrices. By aligning the robot link positions with the corresponding human hand keypoints in pixel space, we ensure that the overlaid robot hand accurately matches the position and orientation of the original hand in the image, achieving precise pixel-level alignment.
III-B Data Quality Evaluation
We evaluate the visual plausibility and semantic consistency of H2R-augmented data using a vision–language similarity metric based on CLIP [42]. The evaluation procedure is illustrated in Figure 2, which also shows six representative pairs of original human frames and their corresponding H2R-augmented robot-centric images.
For each image, we associate a high-level verb–noun action description. Specifically, given an original human frame and its H2R-augmented counterpart, we construct two textual prompts describing the same action from different perspectives: a human-centric prompt (“A human is [action]”) and a robot-centric prompt (“A robotic arm is [action]”). Image–text cosine similarity is then computed using CLIP ViT-B/32 [42], measuring how well each image aligns with its corresponding semantic description.
To quantify data quality at scale, we randomly sample 1,000 image pairs from the pre-training dataset. For each pair, the action phrase is automatically generated using Qwen-2.5-VL [3]. We report the average CLIP similarity between original images and human-centric prompts, and between H2R-augmented images and robot-centric prompts.
The original human images achieve an average similarity score of 28.01, while H2R-augmented images reach 29.83. This result indicates that H2R consistently improves semantic alignment with robot-centric action descriptions. Overall, the CLIP-based evaluation provides a scalable measure of data quality, supporting the effectiveness of H2R in generating visually and semantically coherent training data for robotic visual pre-training.
Figure 2: Illustration of the CLIP-based quality evaluation. We evaluate augmentation quality by comparing CLIP image–text similarity under human-centric and robot-centric action prompts for original versus H2R-augmented images. The results show that H2R improves semantic alignment with robot-centric action descriptions, demonstrating the effectiveness of H2R in generating semantically coherent training data for robotic pre-training.
IV Experiment
TABLE I: Summary of Experimental Settings. This table outlines the hyperparameters, datasets, and configurations used for encoder pre-training, simulation benchmarks, and real-world robotic tasks.
Phase Component Framework Key Configurations
Encoder Pretraining Method MAE ViT-B, 800 epochs, batch size 128, learning rate $4\times 10^{-4}$, $\timesA8$00 GPUs.
R3M ViT-B, 20K steps, batch size 256, learning rate $1\times 10^{-4}$, $\timesA8$00 GPUs.
Dataset SSv2 $\sim$1M frames (subset), original vs. H2R-augmented.
Ego4D 117K clips subset ($\sim$1M frames), original vs. H2R-augmented.
Simulation Setup Benchmark Robomimic 200 epochs, 3 tasks.
RLBench 800 epochs, 3 tasks.
PushT 200 epochs, 1 task.
CortexBench 100 epochs, 3 tasks.
Policy BC 3-layer MLP (Robomimic, CortexBench), default hyperparameters.
DP Diffusion Policy (PushT, RLBench), default hyperparameters.
Real-world Setup Scene UR5-Gripper 3 tasks, keyboard-based teleoperation for collection, 1 camera, 30 demos.
UR5-Leaphand 3 tasks, vision-guided teleoperation for collection, 1 camera, 50 demos.
Dual-arm Franka 3 tasks, homogeneous-arm bilateral teleoperation for collection, 3 cameras, 300 demos.
Dual-arm UR5e 2 tasks, homogeneous-arm bilateral teleoperation for collection, 3 cameras, 300 demos.
Policy DP 300 epochs (UR5-Gripper), 3000 epochs (UR5-Leaphand), 9000 epochs (dual-arm).
ACT 300 epochs (UR5-Gripper), 3000 epochs (UR5-Leaphand), 9000 epochs (dual-arm).
UVA Default UVA configuration, 3000 epochs.
In this section, we conduct a comprehensive experimental evaluation of H2R. Experimental configurations are summarized in Table I. In simulation, we conduct imitation learning experiments across multiple benchmark suites, including studies on generalization across pre-training datasets, comparisons with robotic-data pre-training, and the effect of demonstration density. On real robots, we perform manipulation experiments across multiple platforms and task settings, together with studies on cross-embodiment transfer, compatibility with Vision-Language-Action models, component ablations, and robustness under lighting variations.
IV-A Experimental Setup
Visual Representation Pre-training. We adopt the MAE [22, 53] and R3M [37] frameworks for visual encoder pre-training, both using a Vision Transformer (ViT-Base) [11] architecture. Pre-training is conducted on subsets of SSv2 (1M images) or Ego4D (117K clips).
For each framework, we consider two settings: pre-training on the original human video data and pre-training with additional H2R augmentation. H2R overlays robot embodiments onto human videos using three representative robot types: UR5 with Robotiq Gripper, UR5 with Leaphand, and Franka with Robotiq Gripper. Except for cross-embodiment experiments, the robot embodiment used for H2R augmentation is matched to the robot platform of the downstream task.